the length of a rectangular room is three less than twice its width. Find the dimensions of the perimeter is 54 ft.
Answer by TimothyLamb(4379)

L = 2w - 3
2L + 2w = 54
---
2L + 2w = 54
2(2w - 3) + 2w = 54
4w - 6 + 2w = 54
6w = 60
w = 10
L = 2w - 3
L = 2*10 - 3
L = 23
---
L = 23 ft
w = 10 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php